# FedRecon: Missing Modality Reconstruction in Heterogeneous Distributed Environments

### v2

@@ 5.1. Experimental Setups
Figure 4. Top1-F1 and Top1-accuracy on CrisisMMD under Dedicated FL scenarios, comparing the effects of low ($\alpha=0.1$) and high ($\alpha=5.0$) label imbalance.
Baselines. For evaluating the effectiveness of our MFL framework, we compare FedRecon against a range of both unimodal and multimodal FL methods. Unimodal baselines include FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020b), FedRS (Li and Zhan, 2021), Scaffold (Karimireddy et al., 2020), and FedOpt (Reddi et al., 2020), all implemented with zero-filling for missing modalities following the FedMultimodal (Feng et al., 2023) benchmark. Multimodal baselines include CACMRN (Xiong et al., 2023), FedMMR (Wang et al., 2024), and mmFedMC (Yuan et al., 2024), where zero-filling is likewise adopted in the absence of modality-specific generative modules. We also considered CLAP (Cui et al., 2024), but excluded it in the main comparison due to incompatible settings. To assess the generative performance of our MVAE model, we compare against several representative multimodal generative approaches, including MVAE (Wu and Goodman, 2018), MVTCAE (Hwang et al., 2021), mmJSD (Sutter et al., 2020), MMVAE (Shi et al., 2019), MoPoE-VAE (Sutter et al., 2021b), MMVAE+ (Palumbo et al., 2023), and MMVM (Sutter et al., 2024).
Implementations. For our MFL framework, we implement FedRecon atop the FedMultimodal benchmark (Feng et al., 2023). Following their protocol, we utilize MobileNetV2 (Howard et al., 2017), MobileBERT (Sun et al., 2020), and MFCC (Mel-Frequency Cepstral Coefficients) (Davis and Mermelstein, 1980) to extract features from visual, textual, and audio modalities respectively, enabling efficient and scalable federated training. Therefore, a simple Conv+RNN architecture is used to process visual and audio data, while text is handled by an RNN-only model. For the MELD dataset, which contains only audio and text modalities, we adapt our model structure based on the VAE design provided in a prior implementation [^1] https://github.com/yjlolo/vae-audio to better handle the audio modality. For CrisisMMD, we adopt the vision-language VAE architecture proposed in the original MVAE framework to accommodate the image-text modality pair. We use FedAvg as the baseline method for MELD, while for CrisisMMD we adopt FedOpt. We report Unweighted Average Recall (UAR) on MELD and F1-score on CrisisMMD, along with Top-1 accuracy on both datasets for a more comprehensive evaluation. To evaluate generative baselines, we build upon the official MMVAE+ implementation [^2] https://github.com/epalu/mmvaeplus and refer to their reported metrics for consistency. We also reproduce MMVM [^3] https://github.com/thomassutter/mmvmvae results as an additional reference point. Generative quality is assessed using Fréchet Inception Distance (FID) (Heusel et al., 2017) and generative coherence (Shi et al., 2019) metrics. All experiments are repeated with three different random seeds to ensure robustness and are conducted on a single NVIDIA A100 GPU with 80GB memory. Further implementation details, including those for reproducing closed-source baselines, are provided in Appendix C.
Modality Heterogeneity. We simulate missing modalities by modeling their availability using a Bernoulli distribution, where each modality has an equal probability of being absent. In this work, we adopt the Dedicated FL paradigm (Wang et al., 2024), where both the training and testing sets experience the same missing rate. The data missing rate is varied from 0.1 to 0.9.
Statistical Heterogeneity. For the MELD dataset, we partition the data based on client IDs, as the data is naturally organized by speaker identifiers. In contrast, the CrisisMMD dataset lacks such realistic client partitions, so we use the Dirichlet distribution to partition the data with $\alpha\in\{0.1,5.0\}$, where $\alpha=0.1$ and $\alpha=5.0$ correspond to high and low data heterogeneity, respectively. Figure 7 illustrates the inconsistent data distributions.
5.2. Main Results
Table 1. Comparison of UAR classification results for Dedicated FL on MELD, with $\rho$ varying from 10% to 90%.
@@ 5.2. Main Results
FedRecon 55.44 56.08 54.93 55.47 54.72 53.18 51.34 51.83 49.67
Table 2. Comparison of Top1-accuracy classification results for Dedicated FL on MELD, with $\rho$ varying from 10% to 90%.
@@
FedRecon 69.34 68.94 69.45 69.00 68.97 69.80 67.05 65.61 65.50
MELD Results. The MELD dataset consists of audio and text modalities, which are highly consistent in the information they convey. As a result, the impact of missing modalities on performance is relatively mild, and reconstruction brings limited gains. As shown in Table 1 and Table 2, our method achieves strong performance across different missing ratios. For instance, with 50% modalities missing, FedRecon attains 54.72% UAR, outperforming FedAvg by 1.99%. While the reconstruction process does not lead to significant improvements in classification performance due to inherent modality redundancy, FedRecon still consistently outperforms existing baselines. Notably, under the extreme condition where 90% of modalities are missing, it achieves a Top1-UAR of 49.67% and a Top1-accuracy of 65.50% in the Dedicated FL setting. Additionally, we observe that although the accuracy metric fluctuates noticeably across different random seeds, the trends in UAR and accuracy remain largely consistent. Among the compared MFL methods, both FedMMR and CACMRN incorporate modality reconstruction, which contributes to their superior performance.
@@
Figure 5. Cross-modal generation results on PolyMNIST. Each image is generated from modality m4, showcasing reconstructions and translations into modalities m0–m4.
Figure 6. Cross-modal generation on CUB: images generated from the text modality (m1) to the image modality (m0). Each column corresponds to a set of images generated from the same text input.
CrisisMMD Results. Similarly, due to the sample distribution issue, we evaluate the CrisisMMD dataset using both F1-score and Top1-accuracy. As noted in (Feng et al., 2023), the dataset is particularly sensitive to modality missing. In all missing rate settings, an increase in the missing rate causes a marked decline in performance across all methods. However, occasional performance gains are observed at certain missing ratios, likely because real-world noise in the data can make some modalities less informative—removing them may inadvertently lead to better results. Despite this inherent fluctuation, FedRecon consistently outperforms its counterparts. As evidenced by Figure 4, FedRecon outperforms the unimodal baseline, FedOpt, by an average margin of 2.81% in F1-score under the high imbalance setting ($\alpha=0.1$) across all missing ratios. In the low imbalance setting ($\alpha=5.0$), it maintains an average improvement of 3.84%, further demonstrating its robustness across varying data distributions.
Table 3. Conditional generation performance on PolyMNIST and CUB. Higher coherence and lower FID indicate better performance. We adopt the default hyperparameters used in MMVAE+ for fair comparison.
@@
PolyMNIST and CUB Results. As shown in the Table 3, we directly adopt the reported results of MMVAE+ as a reference benchmark, and further compare FedRecon against MMVM for a comprehensive evaluation. In terms of FID, FedRecon performs relatively poorly on the CUB dataset but yields satisfactory results on PolyMNIST. We speculate that this discrepancy stems from the difference in latent distribution characteristics: the Laplace distribution used for CUB is more challenging to map effectively, whereas the Gaussian distribution adopted for PolyMNIST aligns more naturally with our mapping mechanism. While the reconstruction quality is constrained by this setup, our method achieves sample-level alignment across modalities, which helps improve the overall data distribution and leads to a slight FID improvement over the original MVAE. Regarding generative coherence, since we achieve sample-level alignment, the accuracy of the classifier improves significantly, yielding superior results. FedRecon outperforms both MMVAE+ and MMVM, reaching state-of-the-art performance in this domain. Figure 5 and Figure 6 showcase several randomly selected cross-modal generation results on the PolyMNIST and the CUB test sets.
Figure 7. Plot of the number of samples per class allocated to each client in CrisisMMD, with $\alpha=5.0$ and $\alpha=0.1$.
Figure 8. Ablation results on CrisisMMD under low label imbalance ($\alpha=0.1$).
5.3. Ablation Studies
Contribution of Each Module. We first evaluate the impact of two key modules: the cross-modal alignment reconstruction generator (MMR) and the Global Generator Freezing Strategy (GGFS). To assess their contributions to model performance, we conduct experiments where these modules are excluded from FedRecon. The results underscore the critical role of these components in enhancing overall effectiveness. As shown in Figure 8, FedRecon demonstrates a marked improvement over the baseline, highlighting the necessity of both MMR and GGFS for superior performance. Owing to space limitations, we exclude detailed results on CrisisMMD with $\alpha=5.0$ and the MELD dataset. Nonetheless, we briefly summarize the findings here. On CrisisMMD ($\alpha=5.0$), MMR and GGFS achieve average performance gains of 1.47% and 0.88% respectively across varying missing rates. For MELD, GGFS shows limited or even negative effects, while the MMR module consistently improves UAR by an average of 2.10%.
@@ 5.3. Ablation Studies
Rationale Behind GGFS. We hypothesize that certain clients may hold samples that are inherently difficult to reconstruct due to limitations in local data coverage. This assumption reflects scenarios where local distributions fail to capture the complexity of the global data space. As illustrated in Figure 7, this hypothesis is reasonable and supports the design motivation of GGFS.
5.4. Further Analysis
Figure 9. Examples of Text (m1) generated from Image (m0) on the CUB dataset, using encoder and decoder trained under the MMVAE+ framework.
Figure 10. Examples of Text generated from Image on the CrisisMMD dataset, using the backbone trained with the MMVAE+ framework.
In fact, the mapping mechanism we propose can be applied to any pre-trained VAE. By simply freezing the encoder and decoder, we can train a mapping model to adapt the latent distributions. Essentially, this method aligns the generated data across modalities, ultimately forming a complete data pair. We extend this approach to the architecture trained with MVAE, MMVAE and MMVAE+, and observe that this reconstruction method also facilitates sample alignment to some extent. As illustrated in the Figure 9 abd Figure 10, we present generation results on both the CUB and CrisisMMD datasets using the MMVAE+ framework.
We further evaluate the effectiveness of text-to-image generation on the CrisisMMD dataset. The generated images tend to appear relatively blurry, which may be attributed to the high resolution of the original images, posing challenges for VAE-based synthesis. Given sufficient computational resources, employing diffusion models would likely offer a more suitable alternative.
6. Conclusion
@@
In this paper, we tackle the challenge of missing modalities in the MFL paradigm, which often leads to feature distribution gaps and disrupted inter-modal correlations. We propose FedRecon, a novel framework that incorporates two key components: (1) a cross-modal reconstruction strategy that enables both label-level and sample-level imputation, and (2) a global generator freezing strategy designed to address the issue of unreconstructable samples in Non-IID settings with limited data. Extensive experiments across diverse scenarios demonstrate the effectiveness of FedRecon in enhancing both modality reconstruction and overall performance under Non-IID conditions, paving the way for more robust MFL systems.
Appendix A A Survey on Generative Models
@@
To assess the parameter sizes of various models, we conducted an extensive survey on platforms, including GitHub, Hugging Face, Papers with Code, focusing on representative architectures from three major types of generative models: GANs, VAEs, and Diffusion models. Specifically, we considered ACGAN (Odena et al., 2017), DeGAN (Addepalli et al., 2019), Seg_DeGAN (Bhogale, 2020), MADGAN (Ghosh et al., 2018), MAD-GAN (Li et al., 2019), BEGAN (Berthelot et al., 2017), Boundary-Seeking GAN (Hjelm et al., 2017), StyleGAN2 (Karras et al., 2020) and DragGAN (Pan et al., 2023) from the GAN class, VAE (Kingma et al., 2014), VQ-VAE (van den Oord et al., 2017), VAR (Tian et al., 2024), MVAE (Wu and Goodman, 2018) and MMVAE+ (Palumbo et al., 2023) from the VAE class, as well as DDPM (Ho et al., 2020), DDPM-Seg (Baranchuk et al., 2021), Tab-DDPM (Kotelnikov et al., 2023), Stable Diffusion (Rombach et al., 2022) and Mini-EDM (Karras et al., 2022) from the Diffusion models. These models were selected for their prominence and diversity, and we sampled and computed their parameter sizes to facilitate a comparative analysis.
Table 4. Comparison of Model Parameters and File Sizes
@@
As illustrated in Table 4, for repositories containing multiple model variants, we selected the smallest model for comparison to ensure fairness in parameter and size analysis. VAE exhibits a lightweight advantage over the other two model families, delivering competitive performance while maintaining significantly lower resource consumption. However, VAE also has its limitations. When adapted to multimodal tasks, the quality of the synthesized data often falls short compared to other generative models, exhibiting noticeable artifacts such as blurry edges in reconstructions (Vivekananthan, 2024; Kazerouni et al., 2022). Regarding training time, since each model is designed for different tasks and configurations, a unified evaluation is challenging. Nevertheless, we conducted a comparative test using the most basic implementations of VAE, GAN, and diffusion models on the CIFAR-10 (Krizhevsky et al., 2009) dataset. Among them, VAE demonstrated the fastest training speed.
Appendix B Datasets

### v3

@@ Introduction
Figure 4: Top1-F1 and Top1-accuracy on CrisisMMD under Dedicated FL scenarios, comparing the effects of low ($\alpha=0.1$) and high ($\alpha=5.0$) label imbalance.
@@
CLAP 55.24 55.09 55.34 54.81 53.90 51.84 50.12 48.18 45.18
FedMobile 55.86 55.42 54.83 54.49 54.52 54.26 53.94 52.91 52.08
FedRecon 56.53 56.18 55.95 55.78 55.31 54.90 53.85 53.53 52.84
Table 1: Comparison of UAR classification results for Dedicated FL on MELD, with $\rho$ varying from 10% to 90%.
@@
CLAP 69.71 68.14 67.99 67.86 67.44 65.41 64.51 62.08 60.53
FedMobile 69.55 68.18 67.91 67.45 67.98 67.81 67.10 65.45 64.94
FedRecon 70.83 69.51 69.67 69.16 68.86 68.76 68.13 67.58 66.41
Table 2: Comparison of Top1-acc classification results for Dedicated FL on MELD, with $\rho$ varying from 10% to 90%.
Datasets. To evaluate our MFL framework, we utilize two real-world multimodal datasets: (1) MELD (Poria et al. 2018), a multiparty conversational dataset comprising over 9k utterances with both audio and textual transcripts from the TV series Friends, and (2) CrisisMMD (Alam, Ofli, and Imran 2018), which contains 18.1k tweets accompanied by both visual and textual information, collected during various crisis events. To further assess the generative quality of our MVAE model, we additionally adopt two benchmark multimodal datasets: (3) PolyMNIST (Sutter, Daunhawer, and Vogt 2021a), a synthetic dataset featuring five modalities, where each consists of MNIST digits composited over diverse background images; and (4) CUB (Wah et al. 2011), the Caltech-UCSD Birds dataset, which pairs bird images with fine-grained textual descriptions, posing a challenging setting due to the modality-specific nuances in visual and textual representations.
Baselines. We compare FedRecon against a range of both unimodal and multimodal FL methods. Unimodal baselines include FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020c), FedRS (Li and Zhan 2021), Scaffold (Karimireddy et al. 2020), and FedOpt (Reddi et al. 2020), all implemented with zero-filling for missing modalities following the FedMultimodal (Feng et al. 2023) benchmark. Multimodal baselines include CACMRN (Xiong et al. 2023), FedMMR (Wang et al. 2024), mmFedMC (Yuan et al. 2024), CLAP (Cui et al. 2024) and FedMobile (Liu, Wang, and Yuan 2025) where zero-filling is likewise adopted in the absence of modality-specific generative modules. To assess the generative performance of our MVAE model, we compare against several representative multimodal generative approaches, including MVAE (Wu and Goodman 2018), MVTCAE (Hwang et al. 2021), mmJSD (Sutter, Daunhawer, and Vogt 2020), MMVAE (Shi et al. 2019), MoPoE-VAE (Sutter, Daunhawer, and Vogt 2021b), MMVAE+ (Palumbo, Daunhawer, and Vogt 2023), and MMVM (Sutter et al. 2024).
@@
Figure 5: Cross-modal generation results on PolyMNIST. Each image is generated from modality m4, showcasing reconstructions and translations into modalities m0–m4.
Figure 6: Cross-modal generation on CUB: images generated from the text modality (m1) to the image modality (m0). Each column corresponds to a set of images generated from the same text input.
Implementations. For our MFL framework, we implement FedRecon atop the FedMultimodal benchmark (Feng et al. 2023). We use FedAvg as the baseline method for MELD, while for CrisisMMD we adopt FedOpt. We report Unweighted Average Recall (UAR) on MELD and F1-score on CrisisMMD, along with Top-1 accuracy on both datasets for a more comprehensive evaluation. To evaluate generative baselines, we build upon the official MMVAE+ implementation [^2] https://github.com/epalu/mmvaeplus and refer to their reported metrics for consistency. Generative quality is assessed using Fréchet Inception Distance (FID) (Heusel et al. 2017) and generative coherence (Shi et al. 2019) metrics. All experiments are repeated with three random seeds to ensure robustness and are conducted on a single NVIDIA A100 GPU with 80GB memory. Further implementation details, including those for reproducing closed-source baselines, are provided in Appendix E.
Modality Heterogeneity. We simulate missing modalities by modeling their availability using a Bernoulli distribution, where each modality has an equal probability of being absent. In this work, we adopt the Dedicated FL paradigm (Wang et al. 2024), where both the training and testing sets experience the same missing rate. The data missing rate is varied from 0.1 to 0.9.
Statistical Heterogeneity. For the MELD dataset, we partition the data based on client IDs, as the data is naturally organized by speaker identifiers. In contrast, the CrisisMMD dataset lacks such realistic client partitions, so we use the Dirichlet distribution to partition the data with $\alpha\in\{0.1,5.0\}$, where $\alpha=0.1$ and $\alpha=5.0$ correspond to high and low data heterogeneity, respectively.
Main Results
MELD and CrisisMMD Results. The MELD dataset consists of audio and text modalities, which convey highly consistent information. As a result, the impact of missing modalities on performance is relatively mild, and reconstruction brings limited gains. As shown in Table 1 and Table 2, FedRecon achieves strong performance, e.g., 55.31% UAR at 50% missing modalities, outperforming FedAvg by 2.58%. Notably, even under the extreme condition where 90% of modalities are missing, it achieves a Top1-UAR of 52.84% and a Top1-accuracy of 66.41%. Similarly, due to sample distribution issues, we evaluate CrisisMMD using both F1-score and Top1-accuracy. Performance declines as the missing rate increases across all methods. Despite this inherent fluctuations, FedRecon consistently outperforms its counterparts, achieving an average F1-score improvement of 2.81% over FedOpt at high imbalance ($\alpha=0.1$) and 4.34% under low imbalance ($\alpha=5.0$).
@@
Table 3: Conditional generation performance on PolyMNIST and CUB. Higher coherence and lower FID indicate better performance. We adopt the default hyperparameters used in MMVAE+ for fair comparison.
PolyMNIST and CUB Results. As shown in Table 3, we directly adopt the reported results of MMVAE+ as a reference benchmark, and further compare FedRecon against MMVM for a comprehensive evaluation. In terms of FID, FedRecon performs relatively poorly on the CUB dataset but yields satisfactory results on PolyMNIST. We speculate that this discrepancy stems from the difference in latent distribution characteristics: the Laplace distribution used for CUB is more challenging to map effectively, whereas the Gaussian distribution adopted for PolyMNIST aligns more naturally with our mapping mechanism. While the reconstruction quality is constrained by this setup, our method achieves sample-level alignment across modalities, which helps improve the overall data distribution and leads to a slight FID improvement over the original MVAE. Regarding generative coherence, since we achieve sample-level alignment, the accuracy of the classifier improves significantly, yielding superior results. FedRecon outperforms both MMVAE+ and MMVM, reaching state-of-the-art performance in this domain. Figure 5 and Figure 6 showcase several randomly selected cross-modal generation results on PolyMNIST and CUB test sets. More results can be found in Appendix F.
Figure 7: Ablation results on CrisisMMD under $\alpha=5.0$.
Ablation Studies
Contribution of Each Module. We evaluate the impact of two key modules: the cross-modal alignment reconstruction generator (MMR) and the Global Generator Freezing Strategy (GGFS). To assess their contributions to model performance, we conduct experiments where these modules are excluded from FedRecon. The results underscore the critical role of these components in enhancing overall effectiveness. As shown in Figure 7, FedRecon demonstrates a marked improvement over the baseline, highlighting the necessity of both MMR and GGFS for superior performance. Owing to space limitations, detailed results on CrisisMMD with $\alpha=0.1$ and MELD are provided in Appendix G. Nonetheless, we briefly summarize the findings here. On CrisisMMD ($\alpha=0.1$), MMR and GGFS achieve average F1 gains of 2.79% and 1.55% respectively, across varying missing rates. For MELD, GGFS shows limited impact, with average gains around 0.49%, while the MMR module consistently improves UAR by 2.03% on average.
@@
Rationale Behind GGFS. We hypothesize that certain clients hold samples inherently difficult to reconstruct due to limited local data coverage, reflecting scenarios where local distributions fail to capture global complexity. Supporting evidence for the hypothesis, in the form of label distribution visualizations, along with additional analysis of the GGFS module and qualitative comparisons of generated samples with and without GGFS, are provided in Appendix H.
Further Analysis
Training Efficiency. We provide a detailed training cost analysis in Appendix I, showing that FedRecon remains efficient even with a large number of modalities.
Mapping Complexity Reduction. We can reduce complexity from $O(M^{2})$ to $O(M)$ by assuming each function $T_{n,m}$ is not injective and maps any modality $m$’s latent vectors to $n$’s space. Thus, only one per $T_{n}$ modality is needed. Nevertheless, even with the original $T_{n,m}$, FedRecon trains 80% faster than the full MVAE on PolyMNIST when $|M|=5$. The evaluation of the simplified design, which shows only a very slight performance decrease, is given in Appendix J.
Extensibility to Arbitrary VAE Frameworks. Our mapping mechanism is extensible to any pre-trained VAE by freezing its encoder and decoder and learning to align latent distributions across modalities. We apply it to MMVAE+, with corresponding generation results on CUB presented in Appendix K. Additionally, high-resolution images (e.g., CrisisMMD) pose challenges for VAE-based generation. This can be eased by scaling up the VAE or using diffusion models when resources permit.
Comparison with CLAP. CLAP is the method most similar to ours, but differences in its original setup lead to suboptimal results. Details are discussed in Appendix L.
Conclusion
@@
In this paper, we tackle the challenge of missing modalities in the MFL paradigm, which often leads to feature distribution gaps and disrupted inter-modal correlations. We propose FedRecon, a novel framework that incorporates two key components: (1) a cross-modal reconstruction strategy that enables both label-level and sample-level imputation, and (2) a global generator freezing strategy designed to address the issue of unreconstructable samples in Non-IID settings with limited data. Extensive experiments across diverse scenarios demonstrate the effectiveness of FedRecon in enhancing both modality reconstruction and overall performance under Non-IID conditions, paving the way for more robust and practical MFL systems.
Appendix A Related Work
Multimodal Variational Autoencoder
The Multimodal Variational Autoencoder (MVAE) extends the framework of Variational Autoencoders (VAE) (Kingma, Welling et al. 2014) to enable joint learning and generation across heterogeneous data modalities. Wu and Goodman (Wu and Goodman 2018) pioneered this direction by introducing a scalable architecture capable of learning shared representations in weakly supervised settings, where only a subset of data samples contain all modalities. Subsequent work further advanced the generative coherence and fidelity of MVAEs. Shi et al. (Shi et al. 2019) proposed a mixture-of-experts formulation (MMVAE) that decomposes latent spaces into shared and modality-specific subspaces while ensuring coherent joint generation through multimodal integration. Palumbo et al. (Palumbo, Daunhawer, and Vogt 2023) enhanced this framework by explicitly isolating distributions of shared and private latent factors, thereby improving robustness to variations in latent space design without compromising semantic consistency. Further innovations integrated Diffusion models into the MVAE architecture (Palumbo et al. 2024), demonstrating enhanced unconditional generation performance for complex real-world data modalities. Despite their effectiveness in aligning modalities through joint latent space learning, current methods exhibit limitations in maintaining consistent cross-modal alignment across diverse samples sharing the same label, especially when used on highly variable datasets or data with real-world noise. Notably, FedRecon mitigates these deficiencies through a simple mapping mechanism, demonstrating practical efficacy in cross-modal alignment.
Multimodal Federated Learning
Multimodal Federated Learning (MFL) addresses privacy-preserving collaboration across clients with heterogeneous modality types in real-world applications. Ouyang et al. propose Harmony (Ouyang et al. 2023), which disentangles training into modality-wise federated learning and fusion learning, leveraging balance-aware resource allocation to improve accuracy and speed convergence in Non-IID scenarios. Gao et al. develop FedMVD (Gao et al. 2025), introducing global logit alignment and local angular margin mechanisms to mitigate domain shifts and category imbalance across multimodal nodes. Chen and Zhang present FedMBridge (Chen and Zhang 2024), using a topology-aware hypernetwork to bridge statistical and architectural heterogeneity for personalized multimodal architectures efficiently. While these approaches advance MFL under modality heterogeneity, they generally assume the availability of all modalities on clients. Our method, FedRecon, focuses on missing modality reconstruction in heterogeneous distributed environments, stabilizing training via a frozen global generator replica to reduce distributional shifts caused by client-specific data heterogeneity, thus enhancing usability and robustness in practical scenarios.
Missing Modality Reconstruction in MFL
Reconstructing missing modalities in MFL ensures robust model performance under heterogeneous data distributions. Xiong et al. propose CACMRN (Xiong et al. 2023), a client-adaptive transformer leveraging instance relationships and federated optimization to prevent local overfitting during cross-modal reconstruction. Wang et al. develop FedMMR (Wang et al. 2024), aligning feature spaces via modality synthesis while integrating reconstructed and existing modalities as complementary inputs. Liu et al. design FedMobile (Liu, Wang, and Yuan 2025), prioritizing cross-node latent feature sharing to address extreme modality incompleteness. Saha et al. investigate modality incongruity issues in healthcare FL and explore modality imputation and regularization techniques to reduce missing modality effects (Saha et al. 2025). Yu et al. introduce FedInMM (Yu et al. 2024b), utilizing LSTM-based dynamic weighting and temporal fusion for irregular multimodal data. While effective in reconstructing modalities under limited samples in MFL, these methods overlook sample-level alignment across clients with identical labels. Additionally, they inadequately address instability caused by Non-IID client data during cross-modal generation. FedRecon stabilizes training by freezing a global generator replica to reduce distributional shifts, thereby mitigating fluctuations induced by client-specific data divergence.
Appendix B A Survey on Generative Models
@@
To assess the parameter sizes of various models, we conducted an extensive survey on platforms, including GitHub, Hugging Face, Papers with Code, focusing on representative architectures from three major types of generative models: GANs, VAEs, and Diffusion models. Specifically, we considered ACGAN (Odena, Olah, and Shlens 2017), DeGAN (Addepalli et al. 2019), Seg_DeGAN (Bhogale 2020), MADGAN (Ghosh et al. 2018), MAD-GAN (Li et al. 2019), BEGAN (Berthelot, Schumm, and Metz 2017), Boundary-Seeking GAN (Hjelm et al. 2017), StyleGAN2 (Karras et al. 2020) and DragGAN (Pan et al. 2023) from the GAN class, VAE (Kingma, Welling et al. 2014), VQ-VAE (van den Oord, Vinyals, and kavukcuoglu 2017), VAR (Tian et al. 2024), MVAE (Wu and Goodman 2018) and MMVAE+ (Palumbo, Daunhawer, and Vogt 2023) from the VAE class, as well as DDPM (Ho, Jain, and Abbeel 2020), DDPM-Seg (Baranchuk et al. 2021), Tab-DDPM (Kotelnikov et al. 2023), Stable Diffusion (Rombach et al. 2022) and Mini-EDM (Karras et al. 2022) from the Diffusion models. These models were selected for their prominence and diversity, and we sampled and computed their parameter sizes to facilitate a comparative analysis.
@@
  • 1
    https://github.com/lukedeo/keras-acgan
  • 2
    https://github.com/vcl-iisc/DeGAN
  • 3
    https://github.com/kaushal-py/seg-degan
  • 4
    https://github.com/vinx-2105/MADGAN-DEGAN
  • 5
    https://github.com/Guillem96/madgan-pytorch
  • 6
    https://github.com/eriklindernoren/PyTorch-GAN
  • 7
    https://github.com/eriklindernoren/PyTorch-GAN
  • 8
    https://github.com/XingangPan/DragGAN
  • 9
    https://github.com/NVlabs/stylegan2
  • 10
    https://github.com/bojone/vae
  • 11
    https://github.com/MishaLaskin/vqvae
  • 12
    https://github.com/FoundationVision/VAR
  • 13
    https://github.com/epalu/mmvaeplus
  • 14
    https://github.com/zoubohao/DenoisingDiffusionProbabilityModel-ddpm-
  • 15
    https://github.com/yandex-research/ddpm-segmentation
  • 16
    https://github.com/yandex-research/tab-ddpm
  • 17
    https://github.com/CompVis/stable-diffusion
  • 18
    https://github.com/yuanzhi-zhu/mini˙edm
Table 4: Comparison of Model Parameters and File Sizes
@@
As illustrated in Table 4, for repositories containing multiple model variants, we selected the smallest model for comparison to ensure fairness in parameter and size analysis. VAE exhibits a lightweight advantage over the other two model families, delivering competitive performance while maintaining significantly lower resource consumption. However, VAE also has its limitations. When adapted to multimodal tasks, the quality of the synthesized data often falls short compared to other generative models, exhibiting noticeable artifacts such as blurry edges in reconstructions (Vivekananthan 2024; Kazerouni et al. 2022). Regarding training time, since each model is designed for different tasks and configurations, a unified evaluation is challenging. Nevertheless, we conducted a comparative test using the most basic implementations of VAE, GAN, and diffusion models on the CIFAR-10 (Krizhevsky, Hinton et al. 2009) dataset. Among them, VAE demonstrated the fastest training speed.
Appendix C Pseudocode
The pseudocode of our method is given in Algorithm 1.
Algorithm 1 Two-Stage Federated Training for FedRecon
1:for round $t=1$ to $T$ do
2:  for client $k=1$ to $K$ in parallel do
3:   Stage 1: Local training of MVAE
5:   $\widetilde{T}_{k}^{(t)},T_{k}^{(t)}\leftarrow T^{(t-1)}$
7:   Stage 2: Freeze MVAE and train mapping modules
8:   $\{G_{k,m}^{(t)}\}_{m=1}^{M}\leftarrow\text{Freeze}(\{G_{k,m}^{(t)}\}_{m=1}^{M})$
9:   Update $T_{k}^{(t)}\leftarrow\text{SGD}(T_{k}^{(t)})$ for $L_{2}$ epochs
10:   Stage 3: Update task model
13:  end for
14:  Aggregate:
15:  $\{G_{m}^{(t)}\}_{m=1}^{M}\leftarrow\frac{1}{K}\sum_{k=1}^{K}G_{k,m}^{(t)}$
16:  $T^{(t)}\leftarrow\frac{1}{K}\sum_{k=1}^{K}T_{k}^{(t)}$
17:  $\phi^{(t)}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\phi_{k}^{(t)}$
18:end for
Appendix D Datasets